Two Tests Only.
Professor Davis counts his midterm as 2/3 of the grade, and his final as 1/3 of the grade. Jason scored only 64 on the midterm. What range of scores on the final exam would put Jason’s average between 70 and 79 inclusive

Answer by checkley72(6)

Points for the midterm exam is 2/3 X 64 = 42.67.
Therefore to get a 70 composit score he need to score an additional 27 1/3 points on his final test. But this score only counts for 1/3 of his grade.
For the 70 average he needs an additional 27 1/3 points. But he only gets 1/3 credit therefore the 27 1/3 must be divided by 1/3 = 82. Therefore to get a 70 he needs a final score of at lest 82 on the final test.
For the 79 average he needs an additional 36 1/3 points. But again he only gets 1/3 of credit therefore the 36 1/3 must be divided by 1/3 = 109.
Lets assume he needs a perfect score for the 79 average.